OpenStudy (anonymous):

not sure how to set this up.... Find the volume of the solid enclosed by the paraboloids z=4(x^2+y^2) and z=18−4(x^2+y^2).

OpenStudy (anonymous):

i got as far as r=3/2. not sure what my limits should be or which equation i'm integrating. i looked in the book and there's no example for this in there. really annoying.

OpenStudy (anonymous):

this is supposed to be only using double integrals. not triple. triple is next section.

ganeshie8 (ganeshie8):

Notice that the region of interest is below the paraboloid z = 4(x^2+y^2) and above the paraboloid z = 19-4(x^2+y^2)